OpenStudy (anonymous):

The area of a rectangle is x^2-2x-15 and the length of the rectangle is x+3. Find the width of the rectangle.

OpenStudy (anonymous):

Factor.

OpenStudy (anonymous):

x^2-2x-15 =(x-5)(x+3) L = x+ 3 W = x-5

OpenStudy (anonymous):

ty @surdawi

OpenStudy (anonymous):

\[Area = l*b\]

OpenStudy (anonymous):

Use the Idea...Given....by surdawi

OpenStudy (anonymous):

ok thanks @Yahoo!

OpenStudy (anonymous):

jess please start trying to solve these problems because i will not be there on the test

OpenStudy (anonymous):

Welcome...

OpenStudy (anonymous):

im solving them as i go on my paper and im finding out how i got that answer thank you

OpenStudy (anonymous):

:)